Grade 12th passMechanics

Two soap bubbles of radii r1 and r2 equal to 4cm and 5cm are toiching each other over a common surface s1s2 .Its radii will be

vishal thakur
9 Years agoGrade 12th pass

1 Answer

Ishita Mandal
7 Years ago
Let the radius be R. 
Let surface tension be T
Let pressure outside the bubbles be P
And pressure of the bubbles be P1 and P2 respectively.
Applying the concept of surface tension 
For 1st bubble,
P1-P=4T/r1 .......(1)
For 2nd bubble, 
P2-P=4T/r2 .......(2)
Equation (1)-(2) gives
P1-P2=(4T/r1)-(4T/r2) .......(3)
But we know that P1-P2=Excess pressure which is equal to 4T/R,
Where R is the radius to be calculated as per the question 
So, 
Replacing P1-P2=4T/R in equation (3) we get
4T/R=(4T/r1)-(4T/r2)
1/R=(1/r1)-(1/r2)
1/R=(1/4)-(1/5)
1/R=1/20
R=20cm 
Its radius will be 20 cm